OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (anonymous):

Slope = y2-y1/x2-x1 =-7-3/4-1 =-7 -3/4 +1 -10/5 =-2